OpenStudy (anonymous):

use the following equivalents, along with dimensional analysis, to convert the given unit to the unit indicated. when necessary, round answers to two decimal places. 26 oz to 1lb. (16 oz=1 lb 2000lb = 1 T 1 oz= 28 g 1 lb = 0.45 kg 1 T = 0.9 t)

OpenStudy (anonymous):

26/16=x lbs

OpenStudy (anonymous):

1.63

OpenStudy (anonymous):

How did you get that answer? can you show me the work?

OpenStudy (anonymous):

You have: \[\frac{26oz}{1}*\frac{1 lb}{16oz} \] The ounces divide out and you're left with pounds.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

No problem :)